In an isosceles triangle, two sides are equal. The third side is 75% of the length of one of the equal sides. Find the length of each side when the perimeter is 165cm.
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

In an isosceles triangle, two sides are equal. The third side is 75% of the length of one of the equal sides. Find the length of each side when the perimeter is 165cm.
--------------
Let one of the equal sides be "x"
EQUATION:
x + x + 0.75x = 165
2.75x = 165
x = 60
-----------
Two of the sides are each 60cm
The 3rd side is 0.75*60 = 45cm
===============
Cheers,
Stan H.

Answer by edjones(8007)

Let side a=x b=x c=.75x
Perimeter (P)=a+b+c=165 cm.
So x+x+.75x=165
2.75x=165
divide both sides by 2.75 x=60
Side a=60, b=60, c=.75*60=45
Ed Jones